Ends 10/11 April 2022 midnight
All dates in UTC. Deadlines are 12:00 UTC unless specified.
Overview of all MIB versions and applicability: MIB and FOP
LTP6 Block | STP cycle | IOR delivery opens | IOR deadline | STP POR deadline (on SOC) | Estimated/Planned uplink by MOC ( | MIBs | Start of STP cycle onboard | Comments | SOC ESA responsible | Main operator |
---|---|---|---|---|---|---|---|---|---|---|
1 | 183 |
|
|
| MIB from 18/19 Nov |
| MAINT (SA_REL) Synoptics Operating: All Instruments except PHI. SPICE cal, flat and dark. STIX analysis. | Anik De Groof | usual SCO support | |
1 | 184 |
|
|
|
| MIB from 18/19 Nov |
| Synoptics Operating: All Instruments except PHI (flush only) SPICE cal, flat. Testing in parallel to WOL on first day STIX analysis METIS HSD CLOSING
| David Williams | usual SCO support |
1 | 185 |
|
| 1200utc => feedback on PORs from MOC EOB | (Fri) or weekend? | MIB from 18/19 Nov |
|
EUI (FSI), SPICE, PHI (HRT/FDT), HSD CLOSE
EUI(FSI), SPICE, METIS PHI (HRT/FDT), HSD OPEN Synoptics Operating: All Instruments except PHI. SPICE cal, flat and dark. STIX analysis. | usual SCO support until 23 Dec. DM required, STP issues reviewed as part of DM+ | |
1 | 186 |
|
|
| or , in case of new MIB for 189, earlier date preferred. | MIB from 18/19 Nov |
| TCM type 2 MAINT (any) in NEAR SUN conditions All instruments operating (RS synoptics) PHI helioseismology campaign (= PHI_RAW_FDT_6) spice cal, flat, dark. | David Williams and Jayne Lefort after Dave is on leave | on call support 23 dec -6 jan. from 7 Jan (inc) usual SCO support |
1 | 187 |
|
|
|
| TBC Target MIB for 187 and block 2 (after 14 Jan), ideally possible to apply at SOC from STP 188. |
| SA_ROT All instruments operating (RS synoptics) PHI processing of helioseismology campaign data spice cal, flat, dark. | Anik De Groof | from 7 jan usual sco support |
2 | 188 | / |
|
|
|
| EUI HRI HSD opening at end STP All instruments operating (RS synoptics)
spice cal, flat, dark. PHI raw synoptics run. | |||
2 | 189 |
|
|
|
| SA ROT All instruments operating (RS synoptics) Jitter test during WOL, incl SPICE, PHI and EUI
SPICE HSD CLOSING ahead of LUT upload in next STP | ||||
2 | 190 |
|
|
| TBC Target MIB update for 190 and block 3 (ideally to apply at SOC 191 if possible) |
| TCM type 2 MAINT (any) in NEAR SUN conditions SA Rotation all instruments on (RS synoptics) SPICE HSD opening after LUT update SPICE CE heater set to op limits (i.e., goes cold) spice cal, flat, dark. | Jayne Lefort | ||
3 | 191 | 3/4 February |
|
|
|
| All payload on and operating in RS synoptics or IS science modes. 2 limb-pointing campaigns to East and later West limb for coordinated science with PSP
| |||
3 | 192 | 3/4 February |
| (wed) | (mon) or weekend? | 06:00
|
All Payload in science (and calibration) modes Coordinated science activities (SOOPs):
Calibration activities:
Synoptics at end RSW. (note inferior conjunction does not impose operational constraints)
| |||
3 | 193 | 3/4 February | (tue) | (mon) | (thur) | TBC Target MIB for 193, and block 4 ideally to apply at SOC 194 if possible | | MAINT ANY + TCM Type-2 (in near-sun conditions) SA rotation
SPICE HSD operations for LUT update Calibration activities between RSW1 & 2:
| Anik De Groof | |
4 | 194 | 28 Feb (latest) | (thur) | (wed) | (mon) | 16:30 |
Frequent WOLs. All payload On and operating except Metis Coordinated science activities (SOOPs):
| |||
4 | 195 | 28 Feb (latest) | (fri) or weekend? |
| RSW2+3: VSTP IORs can be planned for and sent, during whole STP.
All payload On and operating. Metis HSD opening for
| |||||
4 | 196 | 28 Feb (latest) |
|
| (fri) or weekend? |
| RSW2+3: VSTP IORs can be planned for and sent, during whole STP. Frequent WOLs. All payload On and operating.
Coordinated science activities (SOOPs):
| |||
44 | 197 | 28 Feb (latest) | (thur) | (wed) | (mon) | Target MIB for next block? if needed available ideally applicable MOC STP 197 and STP SOC 198 |
|
EUI/HRI HS door closure after end RSW3. All payload On and operating. PHI/EUI flatfield calibration, incl 8 offpointing positions Synoptics. SWA must be off at end of STP in preparation for the WOL at start of next LTP The SPICE heater should be powered on during the WOL of 10-Apr-22 12:05–15:05, with the SOU heater being switched back on at 12:35 His off | Jayne Lefort | Mar Sierra |